OpenStudy (anonymous):

How do you find a ratio that is equivalent to 1/6 ?

hero (hero):

Hint multiply 1/6 by any factor of one. For example you can multiply it by 2/2

OpenStudy (anonymous):

So you do 1x6 ?

hero (hero):

You do: \(\huge \frac{1}{6} \times \frac{2}{2} = ?\)

OpenStudy (anonymous):

So I could do that with any number or just 2 ?

hero (hero):

Or you could do \(\huge \frac{1}{6} \times \frac{3}{3} = ?\)

hero (hero):

As long as it is a factor of one, it doesn't matter. Of course any number over itself equals 1.

OpenStudy (anonymous):

Okay, thanks i am stupid !

hero (hero):

You're not stupid. Sometimes it helps for other people to make concepts clearer and easier to understand. Sometimes students have so many things on their mind, it is difficult to focus on what you need to.

OpenStudy (anonymous):

I guess ''/